OpenStudy (ttop0816):

Given the system of equations, what is the x-coordinate of the solution? y = 3 - x 3x + 4y = 1 22 -17/2 -11

OpenStudy (anonymous):

3x+4(3-x)=1

OpenStudy (ttop0816):

sorry! some of the numbers got accidentally erased! Given the system of equations, what is the x-coordinate of the solution? y = 3 - 1/2x 3x + 4y = 1 22 -17/2 -11

OpenStudy (anonymous):

now simplify and solve for ur answer